OpenStudy (firejay5):

Find the inverse of each function. f(x) = -x + 3

OpenStudy (firejay5):

@jim_thompson5910

OpenStudy (anonymous):

how would you solve \[-x+3=5\]

OpenStudy (firejay5):

subtract 3 from both sides

OpenStudy (firejay5):

I got the same function for the inverse and I am not for sure if that's right

jimthompson5910 (jim_thompson5910):

you replace f(x) with y to get y = -x + 3 then you swap x and y: x = -y + 3 and afterward solve for y to get the inverse

jimthompson5910 (jim_thompson5910):

x = -y + 3 x-3 = -y -y = x-3 y = -x+3 so the inverse is the same function f(x) = -x+3

OpenStudy (firejay5):

is that possible

jimthompson5910 (jim_thompson5910):

yes because this function is completely perpendicular to y = x, so it reflects on itself when you reflect it over y = x